Force Dig)--1.102+1.8q-44.3 c. (2.64、$41.39) d.268. S41 .24) MacBook A
Answer #1

Equilibrium point is achieved when:

D(q) = S(q)

-1.1q2 + 1.8q - 44.3 = 3.2q2 + 3q + 2

4.3q2 + 1.2q + 46.3 = 0

There are no roots for this equation, hence there seems to be a problem with this question/typo in the equations. I have tried different permutations and there seems to be a error in equation of D(q). It should be:

D(q) = -1.1q2 + 1.8q + 44.3

When the above equation is taken as D(q), solution is:

D(q) = S(q)

4.3q2 + 1.2q -42.3 = 0

Hence, solving the above equaton, we get q = 3

When q = 3, D(q) = -1.1*9 + 1.8*3 + 44.3 = 39.8

Option: (3, 39.8) is the correct answer